40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Wed May 22 17:48:49 2019, anchal, DailyProgress, TempCtrl, Vacuum Can Temp Sensors restarted with new board Temp-Sensor-Box.jpegBreakoutBox.jpeg
    Reply  Wed Dec 4 17:09:44 2019, anchal, DailyProgress, TempCtrl, Increased range of out-of-loop temperature sensor 
Message ID: 2486     Entry time: Wed Dec 4 17:09:44 2019     In reply to: 2349
Author: anchal 
Type: DailyProgress 
Category: TempCtrl 
Subject: Increased range of out-of-loop temperature sensor 

This has happened few times now that acromag channel for the can heater driver stopped updating according to the PID script and the can gets heated to a very high temperature. This pushes the temperature out of the ranges of the current AD590 temperature sensor board. I have changed the range of channel 2 (this was being used for out-of-loop) to ensure we can still see some meaningful temperature value when such incidents happen. I have replaced R18 from 100k to 27k. The updated table is:

CH No on Board EPICS Channel Name Temperature Conversion Function (ºC) Range (ºC)
1 C3:PSL-TEMP_TABLE = V/0.810875 + 27.30248869 13.860-40.745
2 C3:PSL-TEMP_VACCAN_OOL = V/0.43875 + 33.3115385 8.450-58.155
3 C3:PSL-TEMP_VACCAN_INLOOP = V/1.625 + 33.3115385 26.604-40.020

Weird phenomenon?

  • I'm not sure this problem occurs though. Right now the out-of-loop temperature sensor shows that the temperature of the can is 55 Degree Celsius.
  • This is also not cooling down fast. Last time it took days for it to cool down to set value.
  • The alignment of the cavities change vertically when the can temperature is significantly different from the set value of 34.38 Degree Celcius.
  • This hinders them from locking properly unless I tune the alignment back. But when the can will cool down finally, they will be misaligned again.
  • Also, I just refuse to believe it is actually that hot, but I check the voltage after the transimpedance amplifier in temperature sensor boards and two independent AD590s are reporting this.
  • And anyways, why is the acromag output channel getting frozen anyways.
  • That being said, this is an irreproducible but non-harmful problem yet, so lower priority than the FSS saga going on.
ELOG V3.1.3-